How many cups of a grapefruit must be added to 40 cups of punch that contains 5% grapefruit juice to obtain a punch that is 10% grapefruit?
Answer by Alan3354(69443)

40 cups of 5% --> 95% water = 38 cups of water
That 38 cups will be 90% when the mix is 10% juice
38 = 0.9V
V = 38/0.9 = 380/9 cups of 10% mix
Add 380/9 - 40
= 20/9 cups grapefruit